# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ II.1.1 Construction of Willmore surfaces
Since the Willmore energy is conformally invariant, minimal surfaces of $\mathbb{R}^{3}$, $\mathbb{S}^{3}$ and $\mathbb{H}^{3}$ and their conformal transformations are Willmore. One of the major problems is the understanding of closed Willmore surfaces, which are not conformal transformations of minimal surfaces. In 1985, Pinkall [158] proved that such surfaces exist. Despite the work of Hélein [79], where he constructed the equivalent of Weierstrass–Enneper parametrization in the theory of minimal surfaces, only few methods for constructing explicit examples of Willmore surfaces are known in codimension 1, and most of them are not conformally invariant. For instance, Babich–Bobenko [10] constructed in 1993 examples of Willmore surfaces with a line of umbilic points by gluing minimal tori of $\mathbb{H}^{3}$ along their common boundary in $\partial_{\infty}\mathbb{H}^{3}=\mathbb{R}^{2}$. In 2024, Dall’Acqua–Schätzle [51] proved that any rotationally symmetric Willmore surface with a line of umbilic points is actually a Willmore surface obtained by such construction. To do so, they prove that in this case, the umbilic line should be contained in a plane intersecting the surface orthogonally. By an application of Cauchy–Kovalevskya theorem, they prove that the mean curvature of the parts of the surface above and below this plane, seen as submanifolds of $\mathbb{H}^{3}$, completely vanishes. The $\mathbb{S}^{1}$-equivariant Willmore tori have been classified by Ferus–Pedit [60] (and actually contain the construction of [158]) by analyzing the geodesic flow on these surfaces. The understanding of Willmore graphs has also been subject of numerous works. For example, it is known that entire Willmore graphs with finite energy and entire Willmore graphs of radial functions are flat planes [35, 36, 114]. Simply connected, orientable, complete Willmore surfaces with vanishing Gaussian curvature have been fully classified in [190]. The loop group methods introduced by Hélein [79], has later been developed by Ma [116] and Xia–Shen [191] (see also [54] for a recent survey) and provide a way of constructing Willmore surfaces. However, this is a very abstract method and it would be interesting to know whether one can find more explicit constructions.
@@
An alternative way for such constructions is provided by variational methods, such as minimization or min-max procedures in a given class of surfaces. As an example of application, the Willmore conjecture for genus 1 in codimension 1 has been solved by Marques–Neves [124] in 2013 using a min-max scheme for minimal surfaces combined with Urbano theorem, but remains open for higher genus and general codimension. This conjecture states that up to conformal transformations and in all codimensions, the only minimizer of the Willmore functional in genus 1 is the Clifford torus, whereas in genus larger than 1, the likely candidates are the Lawson surfaces [99, Conjecture 8.4]. Even if there is some numerical evidence [86], its complete resolution remains largely open.
@@
One can ask a few natural questions: can we bound these numbers? Does there exist any immersion realizing these optimization problems? If so, how many are they? Can we classify them? A first issue is that the map $W\colon\mathrm{Imm}(\Sigma,\mathbb{R}^{m})\to\mathbb{R}$ cannot be a Morse function, partly due to the conformal invariance. To overcome this, the third author adopted a viscosity approach in [170], previously introduced for the construction of geodesics in [140]. Instead of considering $W$ alone, he adds a ”smoother” times a small parameter $\sigma>0$ and then let $\sigma\to 0$. He applied successfully this method for a well chosen smoother, in the case $\Sigma=\mathbb{S}^{2}$, and obtained that the values of min-max procedures for $W$ on $\mathbb{S}^{2}$ is the energy of a bubble tree. We refer to [169, 170, 138] for more information about the viscosity method for geodesics or minimal surfaces. A particular case of [170] is the case $m=3$, i.e. the codimension 1 case, which can be applied to the sphere eversion and constitutes a new step towards the $16\pi$-conjecture.
@@
Max–Banchoff [131] proved that any path $\omega\in\Omega$ must contain an immersion $\Phi\in\omega$ having a quadruple point (another proof is given by Hughes [87]). Combining this with Li–Yau inequality shows that $\beta_{0}\geq 16\pi$. As a consequence of [170], we know that the number $\beta_{0}$ is equal to the sum of the energies of finitely many branched Willmore spheres. Hence, the need for a classification of branched Willmore spheres became strongly motivated. As we will explain in the following Section II.1.2, Bryant [28] proved that any smooth Willmore sphere is a conformal transformation of a minimal surface in $\mathbb{R}^{3}$. A first generalization to the branched Willmore spheres has been obtained by Lamm–Nguyen [103] when the total number of branched points counted with multiplicites is at most 3. Michelat and the third author [143] successfully classified all ”variational branched Willmore spheres” which include in particular weak limits of immersed Willmore spheres. The second author in [126] proved that every branched Willmore sphere in ${\mathbb{R}}^{3}$ is the inversion of a minimal surface. This implies in particular that $\beta_{0}$ is a multiple of $4\pi$.
@@
The main result of [138] can be stated as (see for instance [138, Remark 6.5])
@@
In the case where $\Sigma=\mathbb{S}^{2}$ and the ends of $\vec{\Psi}$ are planar (so that $\vec{\Phi}$ as no branch point), Hirsch, Kusner and Mäder-Baumdicker [85, 84], revisiting Michelat’s computations, proved that the Willmore index of $\vec{\Phi}$ is exactly n3n-3, and thus verified Michelat’s conjecture in the case of unbranched Willmore spheres. To do so, they restricted the study of the admissible variations to the following set
@@ II.1.2 Conformal Gauss map approach
One of the major achievement in [28] is the classification of smooth Willmore spheres. Indeed, since there is no holomorphic quartic on $\mathbb{S}^{2}$, every smooth Willmore sphere must satisfy $\mathcal{Q}=0$ and thus, be a conformal transformation of a minimal surface in $\mathbb{R}^{3}$. If $\vec{\Phi}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ is a branched Willmore immersion (for instance obtained in a bubble tree), then the equation $\bar{\partial}\mathcal{Q}=0$ is a-priori not verified at the branch points of $\vec{\Phi}$, hence $\mathcal{Q}$ is a-priori merely meromorphic and it is not clear that Bryant’s classification extends to branched Willmore immersions. Lamm–Nguyen [103] computed an asymptotic expansion of $\mathcal{Q}$ across branch points and proved that $\mathcal{Q}=0$ if the number of branch points of $\vec{\Phi}$ is at most 3, when counted with multiplicities. This idea can also be seen as an application of Liouville theorem for holomorphic functions, as explained in [120, Theorem 2.6.3]. Then Michelat–Rivière [143] proved that a branched Willmore sphere obtained as a weak limit of smooth Willmore surfaces or obtained in a bubbling process of a sequence of smooth Willmore surfaces (called a ”variational branched Willmore sphere”) also verify $\mathcal{Q}=0$. To do so, they compute a refined asymptotic expansion of $\mathcal{Q}$ near a branch point and prove that $\mathcal{Q}$ is actually bounded near branch points where one of the residue of $\vec{\Phi}$ verifies a smallness assumption. They proved in an earlier work [141, 142] that this condition is full-filled for variational branched Willmore spheres, see also [122] for a similar property. The full classification has recently been obtained in [126], by adapting the analysis of the conformal Gauss map developed in [17] in order to obtain $\varepsilon$-regularity on the conformal Gauss map and [127] in order to prove an energy quantization result for Willmore surfaces with bounded Morse index.
The next step, which would be an interesting starting point for the Willmore conjecture, is to understand the geometry of smooth Willmore surfaces $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ when $\Sigma$ has positive genus. On the one hand, Palmer [155] proved that the Bryant’s quartic could be interpreted as the curvature of the conformal Gauss map. Moreover, he proved that if $\mathcal{Q}$ verifies the pointwise condition $|\mathcal{Q}|_{g_{Y}}>1$, then $\Sigma$ is a torus and $\vec{\Phi}$ is unstable. On the other hand, Fisher-Colbrie [61] proved that stable minimal surfaces in positively curved manifold have finite total Gauss curvature. Now that some regularity properties have been understood in [17, 123, 127], it would be intersting to study stability questions from the view-point of the conformal Gauss map.
@@
In higher codimension, the conformal Gauss map has already been studied [148, 56, 117, 115] and there are Willmore spheres which are not conformal transformation of minimal surfaces. For instance, Montiel [148] classified all the Willmore spheres in $\mathbb{R}^{4}$ and proved that in addition to conformal transformations of minimal surfaces, Willmore spheres can also be obtained as a rational curve in $\mathbb{CP}^{3}$ via the Penrose twistor fibration $\mathbb{CP}^{3}\to\mathbb{S}^{4}$. This classification has also been extended in [143] to branched Willmore spheres satisfying some smallness assumption on a residue.
@@
One of the difficulty to such an approach comes from the size of the umbilic set of Willmore surfaces. Contrarily to minimal surfaces, Babich–Bobenko [10] proved that Willmore surfaces can have curves of umbilic points. At these points, the geometry of the conformal Gauss map degenerate. Examples of Willmore surfaces having such curves have been constructed by Babich–Bobenko [10] and by Dall’Acqua–Schätzle [51]. However, the structure of such sets is still not fully understood. Schätzle [177] proved that the umbilic set of a smooth Willmore immersion is composed of isolated points and one dimensional real-analytic manifold without boundary. In the recent work [123], Marque and the second author proved that the structure of the umbilic set provides strong information on the geometry of Willmore immersions and on the value of its energy. For instance if a Willmore surface has a geodesic umbilic line, then it is of Babich–Bobenko type.
@@
What properties does the umbilic lines of closed Willmore surfaces satisfy? Do they satisfy any particular equation? Can we describe their location on the surface? Can we estimate their size in terms of the Willmore energy or any other geometric quantities?
One can also wonder about the geometric interpretation of the Bryant’s quartic. A natural question would be to understand the set of Willmore surfaces having the same given Bryant’s quartic, or even the set of quartics on a given surface that can be achieved as the Bryant’s quartic of some Willmore immersion.
@@
Given two Willmore surfaces $\vec{\Phi}_{1},\vec{\Phi}_{2}\colon\Sigma\to\mathbb{R}^{3}$ such that their Bryant’s quartic satisfies $\mathcal{Q}_{1}=\mathcal{Q}_{2}\neq 0$. Is it possible to relate $\vec{\Phi}_{1}$ to a conformal transformation of $\vec{\Phi}_{2}$? Since the Bryant’s quartic roughly describes the second fundamental form of their conformal Gauss map, the map $\vec{\Phi}_{1}$ should be a conformal transformation of $\vec{\Phi}_{2}$ in the case where the induced metrics of their conformal Gauss map coincide. But it seems wrong if this is not the case. It would be interesting to study the possible differences.
@@
The Poisson problem has been introduced in [47] (analog of the Plateau problem for Willmore surfaces) where the authors solve the Plateau problem for datas given for the immersion and the Gauss map, see also [161]. Due to the conformal invariance of the Willmore energy, it would be natural to solve the Plateau problem for conformally invariant data on the boundary.
@@
What is the optimal regularity assumption on the given boundary datas in order to solve the Poisson problem? How many solutions are there?
@@ II.1.3 Willmore flow
Geometric flows are designed to produce explicit paths leading to critical points of a given functional. They are useful both from a theoretic and from a practical view-point. For instance, Conjecture II.3 is an example of application of the Willmore flow, which roughly says that there is no smooth Willmore sphere of energy strictly between the energy of $4\pi$ and $16\pi$. One could also think about the Willmore flow as a first step before studying the Canham–Helfrich flow modelling the evolution of membranes of the cells, see for instance [25, 173] and the references therein.
The main analytical difficulty of the Willmore flow (despite the fact that it is a geometric flow, and thus has to be degenerate) is the fact that it is of order 4, meaning that there exists no comparison principle in the spirit of the mean curvature flow. For instance, if an initial data is included in the interior of a given sphere $S$, one can a-priori not say that the flow remains inside the interior of $S$. Geometric flows of order 4 have original been explored in the context of biharmonic surfaces [57]. The Willmore flow has first been introduced in 2001 independentely by Kuwert–Schätzle [101] and Simonett [181]. They proved the existence of a solution in short time, and proved that if the initial data has small energy, then the flow smoothly converges to the round sphere. In 2003, Mayer–Simonett [132] proved that the Willmore flow does not preserve embeddedness of the initial data. Blatt [22] constructed in 2009 an example of initial data of spherical type for which a singularity occurs in the Willmore flow. More recently, Dall’Acqua–Müller–Schätzle–Spener [50] produced examples of flows made of tori of revolutions, and also produced degenerating examples. This flow is still widely unexplored. The parametric approach has been introduced by Palmurella–Rivière [156, 157], where the authors show that there are only two possible kind of singularities in finite time: either a concentration of energy with a possible branch point, or a degeneration of the conformal class.
@@
In order to study the possible degeneracies, one can rely on the existing literature on the bubbling analysis of sequences $\vec{\Phi}_{k}\colon\Sigma\to\mathbb{R}^{3}$ of Willmore surfaces. When the underlying complex structure of $(\Sigma,g_{\vec{\Phi}_{k}})$ remains bounded in the moduli space, Bernard–Rivière [19] proved in 2014 that some bubbles can appear, but the energy is always preserved. Then Laurain–Rivière [106] proved in 2018 that when the underlying complex structure degenerates, then only one residue is responsible for a possible loss of energy. This has recently been extended in [112], where Li–Yin–Zhou prove that this residue precisely describes the loss of energy. In 2021, Marque [122] proved that the set of Willmore immersions with energy less than or equal to $12\pi$ is compact. In 2025, the second author [127] proved that the energy is preserved if $\vec{\Phi}_{k}$ has bounded Morse index, independently of the underlying complex structure. Recently, Michelat [139] proved the lower semi-continuity of the Morse index under some smallness assumption of the residue (which corresponds to the assumption in [106]). Since the analysis of Willmore immersions is now understood, it should be possible to study the case of the flow.
@@
As an application, one can think about the $16\pi$-conjecture. In a recent work [118], Mäder-Baumdicker–Seidel studied the possible singularities that occur for an initial data $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ with Willmore energy strictly less than $16\pi$. They proved that if $W(\vec{\Phi}_{0})\leq 12\pi$, then either the flow starting from $\vec{\Phi}_{0}$ converges to a round sphere or the flow converges to two round spheres glued by a catenoid neck. If $W(\vec{\Phi}_{0})<16\pi$, then other singularities might arise.
Open Question II.18.
Given a immersion $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{m}$ satisfying $W(\vec{\Phi}_{0})<16\pi$, under which conditions the Willmore flow $(\vec{\Phi}_{t})_{t>0}$ starting from $\vec{\Phi}_{0}$ does not develop any singularity as long as $12\pi<W(\vec{\Phi}_{t})<16\pi$?

### v3

@@ II.1.1 Construction of Willmore surfaces
Since the Willmore energy is conformally invariant, minimal surfaces of $\mathbb{R}^{3}$, $\mathbb{S}^{3}$ and $\mathbb{H}^{3}$ and their conformal transformations are Willmore. One of the major problems is the understanding of closed Willmore surfaces, which are not conformal transformations of minimal surfaces. In 1985, Pinkall [161] proved that such surfaces exist. Despite the work of Hélein [80], where he constructed the equivalent of Weierstrass–Enneper parametrization in the theory of minimal surfaces, only few methods for constructing explicit examples of Willmore surfaces are known in codimension 1, and most of them are not conformally invariant. For instance, Babich–Bobenko [10] constructed in 1993 examples of Willmore surfaces with a line of umbilic points by gluing minimal tori of $\mathbb{H}^{3}$ along their common boundary in $\partial_{\infty}\mathbb{H}^{3}=\mathbb{R}^{2}$. In 2024, Dall’Acqua–Schätzle [51] proved that any rotationally symmetric Willmore surface with a line of umbilic points is actually a Willmore surface obtained by such construction. To do so, they prove that in this case, the umbilic line should be contained in a plane intersecting the surface orthogonally. By an application of Cauchy–Kovalevskaya theorem, they prove that the mean curvature of the parts of the surface above and below this plane, seen as submanifolds of $\mathbb{H}^{3}$, completely vanishes. The $\mathbb{S}^{1}$-equivariant Willmore tori have been classified by Ferus–Pedit [60] (and actually contains the construction of [161]) by analyzing the geodesic flow on these surfaces. The understanding of Willmore graphs has also been subject of numerous works. For example, it is known that entire Willmore graphs with finite energy and entire Willmore graphs of radial functions are flat planes [35, 36, 117]. Simply connected, orientable, complete Willmore surfaces with vanishing Gaussian curvature have been fully classified in [194]. The loop group methods introduced by Hélein [80], has later been developed by Ma [120] and Xia–Shen [195] (see also [54] for a recent survey) and provide a way of constructing Willmore surfaces. However, this is a very abstract method and it would be interesting to know whether one can find more explicit constructions.
@@
An alternative way for such constructions is provided by variational methods, such as minimization or min-max procedures in a given class of surfaces. As an example of application, the Willmore conjecture for genus 1 in codimension 1 has been solved by Marques–Neves [128] in 2012 using a min-max scheme for minimal surfaces combined with Urbano theorem, but remains open for higher genus and general codimension. This conjecture states that up to conformal transformations and in all codimensions, the only minimizer of the Willmore functional in genus 1 is the Clifford torus, whereas in genus larger than 1, the likely candidates are the Lawson surfaces [100, Conjecture 8.4]. Even if there is some numerical evidence [87], its complete resolution remains largely open. The case of surfaces enjoying a priori the same symmetries as the Lawson surfaces has been recently studied in [101].
@@
One can ask a few natural questions: can we bound these numbers? Does there exist any immersion realizing these optimization problems? If so, how many are they? Can we classify them? A first issue is that the map $W\colon\mathrm{Imm}(\Sigma,\mathbb{R}^{m})\to\mathbb{R}$ cannot be a Morse function, partly due to the conformal invariance. To overcome this, the third author adopted a viscosity approach in [173], previously introduced for the construction of geodesics in [144]. Instead of considering $W$ alone, he adds a ”smoother” times a small parameter $\sigma>0$ and then let $\sigma\to 0$. He applied successfully this method for a well chosen smoother, in the case $\Sigma=\mathbb{S}^{2}$, and obtained that the values of min-max procedures for $W$ on $\mathbb{S}^{2}$ is the energy of a bubble tree. We refer to [172, 173, 142] for more information about the viscosity method for geodesics or minimal surfaces. A particular case of [173] is the case $m=3$, i.e. the codimension 1 case, which can be applied to the sphere eversion and constitutes a new step towards the $16\pi$-conjecture, originally formulated by Kusner in the 1980s (see for instance [102]).
@@
Max–Banchoff [135] proved that any path $\omega\in\Omega$ must contain an immersion $\Phi\in\omega$ having a quadruple point (another proof is given by Hughes [88]). Combining this with Li–Yau inequality shows that $\beta_{0}\geq 16\pi$. As a consequence of [173], we know that the number $\beta_{0}$ is equal to the sum of the energies of finitely many branched Willmore spheres. In [146], Michelat and the third author proved an energy quantization of the Gauss map in the $W^{1,(2,1)}$ topology for sequences of immersions arising in the viscosity method. Hence, the need for a classification of branched Willmore spheres became strongly motivated. As we will explain in the following Section II.1.2, Bryant [28] proved that any smooth Willmore sphere is a conformal transformation of a minimal surface in $\mathbb{R}^{3}$. A first generalization to the branched Willmore spheres has been obtained by Lamm–Nguyen [106] when the total number of branched points counted with multiplicities is at most 3. Michelat and the third author [145] successfully classified all ”variational branched Willmore spheres” which include in particular weak limits of immersed Willmore spheres. The second author in [130] proved that every branched Willmore sphere in ${\mathbb{R}}^{3}$ is the inversion of a minimal surface. This implies in particular that $\beta_{0}$ is a multiple of $4\pi$.
@@
The main result of [142] can be stated as (see for instance [142, Remark 6.5])
@@
In the case where $\Sigma=\mathbb{S}^{2}$ and the ends of $\vec{\Psi}$ are planar (so that $\vec{\Phi}$ as no branch point), Hirsch, Kusner and Mäder-Baumdicker [86, 85], revisiting Michelat’s computations, proved that the Willmore index of $\vec{\Phi}$ is exactly n3n-3, and thus verified Michelat’s conjecture in the case of unbranched Willmore spheres. To do so, they restricted themselves to the study of the following admissible variations
@@ II.1.2 Conformal Gauss map approach
One of the major achievement in [28] is the classification of smooth Willmore spheres. Indeed, since there are no non-zero holomorphic quartic differentials on $\mathbb{S}^{2}$, every smooth Willmore sphere must satisfy $\mathcal{Q}=0$ and thus, be a conformal transformation of a minimal surface in $\mathbb{R}^{3}$. If $\vec{\Phi}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ is a branched Willmore immersion (for instance obtained in a bubble tree), then the equation $\bar{\partial}\mathcal{Q}=0$ is a-priori not verified at the branch points of $\vec{\Phi}$, hence $\mathcal{Q}$ is a-priori merely meromorphic and it is not clear that Bryant’s classification extends to branched Willmore immersions. Lamm–Nguyen [106] computed an asymptotic expansion of $\mathcal{Q}$ across branch points and proved that $\mathcal{Q}=0$ if the number of branch points of $\vec{\Phi}$ is at most 3, when counted with multiplicities. This idea can also be seen as an application of Liouville theorem for holomorphic functions, as explained in [124, Theorem 2.6.3]. Then Michelat–Rivière [145] proved that a branched Willmore sphere obtained as a weak limit of smooth Willmore surfaces or obtained in a bubbling process of a sequence of smooth Willmore surfaces (called a ”variational branched Willmore sphere”) also verifies $\mathcal{Q}=0$, by showing that the limit of Bryant’s quartics must remains holomorphic. The full classification has recently been obtained by the second author in [130], by adapting the analysis of the conformal Gauss map developed in [17] in order to obtain $\varepsilon$-regularity on the conformal Gauss map and [131] in order to prove an energy quantization result for Willmore surfaces with bounded Morse index.
The next step, which would be an interesting starting point for the Willmore conjecture, is to understand the geometry of smooth Willmore surfaces $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ when $\Sigma$ has a positive genus. On the one hand, Palmer [158] proved that Bryant’s quartic can be interpreted as the curvature of the conformal Gauss map. Moreover, he proved that if $\mathcal{Q}$ verifies the pointwise condition $|\mathcal{Q}|_{g_{Y}}>1$, then $\Sigma$ is a torus and $\vec{\Phi}$ is unstable. On the other hand, Fisher-Colbrie [61] proved that stable minimal surfaces in a positively curved manifold have finite total Gauss curvature. Now that some regularity properties have been understood in [17, 127, 131], it would be interesting to study stability questions from the viewpoint of the conformal Gauss map.
@@
In higher codimension, the conformal Gauss map has already been studied [151, 56, 121, 119] and there are Willmore spheres that are not conformal transformations of minimal surfaces. For instance, Montiel [151] classified all the Willmore spheres in $\mathbb{R}^{4}$ and proved that in addition to conformal transformations of minimal surfaces, Willmore spheres can also be obtained as a rational curve in $\mathbb{CP}^{3}$ via the Penrose twistor fibration $\mathbb{CP}^{3}\to\mathbb{S}^{4}$. This classification has also been extended in [145] to branched Willmore spheres arising as limits of smooth Willmore surfaces. The classification of smooth Willmore spheres in arbitrary codimension is much more involved and has recently been studied in [118].
@@
One of the difficulties with such an approach comes from the size of the umbilic set of Willmore surfaces. Unlike minimal surfaces, Babich–Bobenko [10] proved that Willmore surfaces can have curves of umbilic points. At these points, the geometry of the conformal Gauss map degenerates. Examples of Willmore surfaces having such curves have been constructed by Babich–Bobenko [10] and by Dall’Acqua–Schätzle [51]. However, the structure of such sets is still not fully understood. Schätzle [180] proved that the umbilic set of a smooth Willmore immersion is composed of isolated points and one dimensional real-analytic manifolds without boundary. In the recent work [127], Marque and the second author proved that the structure of the umbilic set provides strong information on the geometry of Willmore immersions and on the value of its energy. For instance, if a Willmore surface has a geodesic umbilic line, then it is of Babich–Bobenko type.
@@
What properties do the umbilic lines of closed Willmore surfaces satisfy? Do they satisfy any particular equation? Can we describe their location on the surface? Can we estimate their size in terms of the Willmore energy or any other geometric quantities?
One can also wonder about the geometric interpretation of the Bryant’s quartic. A natural question would be to understand the set of Willmore surfaces having the same given Bryant’s quartic, or even the set of quartics on a given surface that can be achieved as the Bryant’s quartic of some Willmore immersion. For instance, we can consider two immersions $\vec{\Phi}_{1},\vec{\Phi}_{2}\colon\Sigma\to\mathbb{R}^{3}$ having the same Bryant’s quartic. Then one can roughly think of this restriction as a relation between the second fundamental forms of their conformal Gauss maps. If their induced metrics $g_{\vec{\Phi}_{1}}$ and $g_{\vec{\Phi}_{2}}$ are also conformal, then one can hope that these two conformal Gauss maps coincide, and thus $\vec{\Phi}_{1}$ and $\vec{\Phi}_{2}$ should be a conformal transformation of the other one.
@@
Given two Willmore surfaces $\vec{\Phi}_{1},\vec{\Phi}_{2}\colon\Sigma\to\mathbb{R}^{3}$ such that their Bryant’s quartic satisfies $\mathcal{Q}_{1}=\mathcal{Q}_{2}\neq 0$. Is it possible to relate $\vec{\Phi}_{1}$ to a conformal transformation of $\vec{\Phi}_{2}$? It would be interesting to study the possible differences.
@@
The Poisson problem has been introduced in [47] (analog of the Plateau problem for Willmore surfaces), where the authors solve the Plateau problem for data given for the immersion and the Gauss map, see also [164]. Due to the conformal invariance of the Willmore energy, it would be natural to solve the Plateau problem for conformally invariant data on the boundary.
@@
What is the optimal regularity assumption on the given boundary data in order to solve the Poisson problem? How many solutions are there?
@@ II.1.3 Willmore flow
Geometric flows are designed to produce explicit paths leading to critical points of a given functional. They are useful both from a theoretic and from a practical viewpoint. For instance, Conjecture II.3 is an example of application of the Willmore flow, which roughly says that there is no smooth Willmore sphere of energy strictly between the energy of $4\pi$ and $16\pi$. One could also think about the Willmore flow as a first step before studying the Canham–Helfrich flow modeling the evolution of membranes of the cells, see for instance [25, 176] and the references therein.
The main analytical difficulty of the Willmore flow (despite the fact that it is a geometric flow, and thus has to be degenerate) is the fact that it is of order 4, meaning that there exists no comparison principle in the spirit of the mean curvature flow. For instance, if an initial data is included in the interior of a given sphere $S$, one can a-priori not say that the flow remains inside the interior of $S$. Geometric flows of order 4 have originally been introduced in the context of biharmonic surfaces [57]. The Willmore flow was first introduced in 2001 independently by Kuwert–Schätzle [104] and Simonett [184]. They proved the existence of a solution in short time, and proved that if the initial data has small energy, then the flow smoothly converges to the round sphere. In 2003, Mayer–Simonett [136] proved that the Willmore flow does not preserve the embeddedness of the initial data. Blatt [22] constructed in 2009 an example of initial data of spherical type for which a singularity occurs in the Willmore flow. More recently, Dall’Acqua–Müller–Schätzle–Spener [50] produced examples of flows made of tori of revolution, and also produced degenerating examples. This flow is still widely unexplored. The parametric approach has been introduced by Palmurella–Rivière [159, 160], where the authors show that there are only two possible kinds of singularity in finite time: either a concentration of energy with a possible branch point or a degeneration of the conformal class.
@@
In order to study the possible degeneracies, one can rely on the existing literature on the bubbling analysis of sequences $\vec{\Phi}_{k}\colon\Sigma\to\mathbb{R}^{3}$ of Willmore surfaces. When the underlying complex structure of $(\Sigma,g_{\vec{\Phi}_{k}})$ remains bounded in the moduli space, Bernard–Rivière [19] proved in 2014 that some bubbles can appear, but the energy is always preserved. Then Laurain–Rivière [109] proved in 2018 that when the underlying complex structure degenerates, then only one residue is responsible for a possible loss of energy. This has recently been extended in [115], where Li–Yin–Zhou proved that this residue precisely describes the loss of energy. In 2021, Marque [126] proved that the set of Willmore immersions with energy less than or equal to $12\pi$ is compact. In 2025, the second author [131] proved that the energy is preserved if $\vec{\Phi}_{k}$ has bounded Morse index, independently of the underlying complex structure. Recently, Michelat [143] proved the lower semi-continuity of the Morse index under some smallness assumption of the residue (which corresponds to the assumption in [109]). Since the analysis of Willmore immersions is now understood, it should be possible to study the case of the flow.
@@
As an application, one can think about the $16\pi$-conjecture and the classification of singularities along the Willmore flow. Lamm–Nguyen [106] proved that if we start the flow with an initial data $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ verifying $W(\vec{\Phi}_{0})\leq 12\pi$, then the flow starting from $\vec{\Phi}_{0}$ can only develop catenoid-type singularities. In a recent work [122], Mäder-Baumdicker–Seidel identified 4 distinct homotopy classes in the set of immersed spheres with Willmore energy less than or equal to $12\pi$. Two of them contain round spheres and the two others lead to a notion of unavoidable singularities along the flow. A partial study of the set of immersions with energy less than $16\pi$ is also studied in [122].
Open Question II.18 (Question 1.13 in [122]).
Is it possible to classify the homotopy classes of the set of immersions with Willmore energy less than $16\pi$? In particular, given an immersion $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ satisfying $W(\vec{\Phi}_{0})<16\pi$, under which conditions the Willmore flow $(\vec{\Phi}_{t})_{t}$ starting from $\vec{\Phi}_{0}$ lies in the homotopy class of a round sphere when $W(\vec{\Phi}_{t})<12\pi$?